lmiddlebusher:

a first order reaction has a 1/2 life of 63.4 mins at 225 degrees C. Calculate the rate constant for the reaction

lmiddlebusher:

I just need help with the equation

Elite:

Well to find the half life, you'd use \[t1/2=\frac{ \ln (2)}{ k }\] where "k" is the rate constant. So just flip the rate constant and half life around: \[k=\frac{ \ln (2) }{ t1/2 }\] Divide the natural log of 2 by the half life (63.4 minutes) and you end up with the answer in seconds.

fwval:

e